OpenStudy (anonymous):

zack places 72 sweets of chocolate,orange and manggo flavours on a table. 19 sweets are of chocolate flavour. a sweet is picked at random from the table. the probability of picking an orange-flavoured sweet is 7/18. how many mango-flavoured sweets are there on the table.

OpenStudy (anonymous):

u need to find out how many mango flavored sweets are there. so in total thereare 72 sweets. u kno how many choco u have. n u have a probability for amount of orange. so lets find amount of orange

OpenStudy (anonymous):

7/18=x/72 and swolve for x

OpenStudy (anonymous):

cross multiply and u get x=28

OpenStudy (anonymous):

so amount of mango is total amount of sweets-amount of orange-amount of choco mango=72-19-28 mango=25 so probability of getting mango is: 25/72